OpenStudy (anonymous):

A six-sided number cube labeled 1 through 6 is rolled 500 times. An odd number is rolled 325 times. Compare the experimental probability of rolling an odd number with the theoretical probability of rolling an odd number and select one of the statements below that best describes the situation. The experimental probability and theoretical probability are the same. The experimental probability is larger than the theoretical probability. The experimental probability is smaller than the theoretical probability. There is not enough information to determine the relative frequency.

OpenStudy (kirbykirby):

Well the experimental probability is (# of successful trials)/(# all trials) = 325/500

OpenStudy (kirbykirby):

For the theoretical probability, notice that rolling an odd number means you can roll either 1, 3 or 5. So, the theoretical probability of rolling an odd number is: \[ {P\{ 1,3,5\}}=\frac{3}{6}\]

OpenStudy (anonymous):

so would it be B

OpenStudy (kirbykirby):

yes =]